A weak acid has a dissociation constant of Ka = 2.50×10−2.

A. Calculate the degree of dissociation for a 9.7×10−2 m solution of this acid using the Debye-Huckel limiting law.

B. Calculate the degree of dissociation for a 9.7×10−2 m solution of this acid that is also 0.200 m in KCl from the Debye-Huckel limiting law using an iterative calculation until the answer is constant in the second decimal place.

C. Repeat the calculation in (B) using the value of γ± = 0.718 for KCl.
Answer #1

Ans: Given Dissociation constant of weak acid, Kas 215x16. s) ka= 0.025.4 A Given - 2 . of solution. dissociation for 9.9x10I (2) = 0.03% motika In Yt = -1•113x1 x 10.03% 1 : -0.2295 = 20-2295 =) = 0.1949 when f = 0.1949 them from eq 6 . m” (0-1949)-> 9 = 0.781 when vi = 018 them from 090 m (0.609) (0-097-m) = m +0.04m - (3:28x10%) = 0 ME -0.04 + 160-045 – 4(1)(-(3-88x16Bikel Z=1 7.31 Given 1:0.2m Activity woefficient is I= 0;2 (2) = 0.2. mellka mY = --173 x 1 x Jo-2 = -0.524 ě0.524 CalculateI: 0.2-0.0145 -I = 0-245 mol kg. Y : -1.173X1 X Jo.245 -0.58 -) = 20:58 -) 11 = 0-559 Recalculate r , from M (0.559) eq o =S. Given (t = 0.718 from eq m(0-418) 0.091-m mt 0.04 m - (3088x103) = 0 -0.04 + Jo.04)=-4 (1) 6-(3-88X163)) me - 20.025 i p